OpenStudy (anonymous):

5sqrt6x+8sqrt14x-9sqrt6x+18sqrt14x

OpenStudy (anonymous):

i dont know how to solve this one....

OpenStudy (anonymous):

Like this..\[=5\sqrt{6x}+8\sqrt{14x}-9\sqrt{6x}+18\sqrt{14x}\]First solve to get \[=-4\sqrt{6x}+26\sqrt{14x}\]Now square this \[=(-4\sqrt{6x}+26\sqrt{14x})^2\]

OpenStudy (anonymous):

so thats not the answer?

OpenStudy (anonymous):

@experimentX

OpenStudy (jiteshmeghwal9):

now for solving thi further use\[\huge{(a+b)^2=a^2+b^2+2ab}\]here, a=\(-4\sqrt{6x}\) b=\(26\sqrt{14x}\)

OpenStudy (jiteshmeghwal9):

ohh ! @Miyuru plz tell me why u square it ??

OpenStudy (anonymous):

Well I thought that would simplify further... :D

OpenStudy (jiteshmeghwal9):

@kaleighdunn u only need to combine like terms\[5\sqrt{6x}+8\sqrt{14x}-9\sqrt{6x}+18\sqrt{14x}\]\[5\sqrt{6x}-9\sqrt{6x}+18\sqrt{14x}+8\sqrt{14x}\]\[-4\sqrt{6x}+26\sqrt{14x}\]

OpenStudy (jiteshmeghwal9):

that's ur answer :)

OpenStudy (anonymous):

I think I went too far.. :D

OpenStudy (jiteshmeghwal9):

;)

OpenStudy (anonymous):

thank you so much!

OpenStudy (jiteshmeghwal9):

yw :)